### v1

# Triad: Vision Foundation Model for 3D Magnetic Resonance Imaging

Abstract
Vision foundation models (VFMs) are pre-trained on extensive image datasets to learn general representations for diverse types of data. These models can subsequently be fine-tuned for specific downstream tasks, significantly boosting performance across a broad range of applications. However, existing vision foundation models that claim to be applicable to various radiology tasks are mostly pre-trained on 3D computed tomography (CT), which benefits from the availability of extensive 3D CT databases. Significant differences between CT and magnetic resonance imaging (MRI) in imaging principles, signal characteristics, and data distribution may hinder their practical performance and versatility in MRI-specific applications. Here, we propose Triad, a vision foundation model for 3D MRI. Triad adopts a widely used autoencoder architecture to learn robust representations from 131,170 3D MRI volumes and uses organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. The above pre-training dataset is called Triad-131K, which is currently the largest 3D MRI pre-training dataset. We evaluate Triad across three tasks, namely, organ/tumor segmentation, organ/cancer classification, and medical image registration, in two data modalities (within-domain and out-of-domain) settings using 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 6.88% compared to nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across two datasets. Our study demonstrates that pre-training can maximize performance when the data modalities and organs of upstream and downstream tasks are consistent. This work highlights the value of large-scale pre-training techniques for downstream tasks in 3D MRI. By open-sourcing Triad’s weights, code, and data, we aim to enhance the adaptability and reliability of AI solutions for 3D MRI in radiology.
1 Introduction
Each year, over 40 million magnetic resonance imaging (MRI) scans are performed in the United States, an average of 107.5 scans per 1,000 persons oecd_mri_units_2023 ; shah2023review . Globally, the annual total of MRI scans ranges between 100 to 150 million mri_number_worldwide . This has led to a growing demand for automated analysis tools https://doi.org/10.1002/mp.16844 . In recent years, Foundation Model (FM)-driven image analysis has shown significant advancements. However, these foundation models have primarily been tailored for general computer vision tasks and are trained on numerous natural image datasets to acquire general representations applicable to a wide range of data Caron_2021_ICCV ; he2022masked ; oquab2023dinov2 . They can then be fine-tuned for various specific downstream tasks, leading to significant enhancements in performance across different applications. This paradigm shift has also been widely adopted in clinical modalities (including 2D and 3D data), which has demonstrated notable improvements moor2023foundation ; chen2024towards ; Zhao2024 ; Safari_2024 .
Figure 1: Overview of Triad training and evaluation. a. Triad pre-training strategy. Triad implements the reconstruction task based on autoencoders and uses L1 loss for optimization. Imaging descriptions are embedded into vector space to form a distribution, which serves as a supervisory signal to constrain the distribution of visual modalities using Log-ratio loss kim2019deep . The two losses are optimized simultaneously in a multi-task manner. Triad is then evaluated in within-domain tasks and out-of-domain tasks. These include within-domain 3D MRI segmentation, classification, and registration tasks (tasks b, c, and d). And unseen 3D CT/MRI segmentation, classification, and registration tasks (tasks e, f, and g).
Figure 2: An overview of the Triad-131K pre-training dataset. a. Describes the name and scale distribution of each dataset in Triad-131K. b. We compare the parameter scale and data scale used by Triad and existing foundation models, and it is obvious that Triad crushes existing models on both scales. c. Shows examples of visual volumetric modality and textual modality in Triad-131K. d. Shows the dataset scale distribution of three organs: brain, breast, and prostate.
Unfortunately, the potential of foundational models specifically for 3D MRI remains largely unexplored. There are two key limitations: Firstly, although previous general medical foundation models assert the capability to generalize to 3D MRI, substantial differences in imaging principles, signal characteristics, and data distribution between other modalities and MRI may hinder their practical performance and generalizability in MRI-specific applications perez2021torchio ; cardoso2022monai ; https://doi.org/10.1002/mp.17675 . For instance, notable general medical foundation models wu2024voco ; li2024well ; tang2022self have predominantly been pre-trained on 3D computed tomography (CT). Although some foundation models leverage mixed modalities during the pre-training phase—including MRI, CT, positron emission tomography, and microscopy ye2024continual ; gao2024training ; isensee2021nnu —there exists a substantial data imbalance. For example, MedCoSS ye2024continual relied on only 3,000 MRI scans, compared to 10,000 CT scans. Secondly, while there have been attempts to develop 3D MRI-specific foundation models, these efforts have typically focused on data from a single organ cox2024brainsegfounder ; kim2023empirical , primarily emphasizing on T1- (T1-w) and T2-weighted (T2-w) images while overlooking the significance of other MRI sequences. For instance, Brainsegfounder cox2024brainsegfounder uses brain T1-w and T2-w images from approximately 80,000 healthy subjects for pre-training, followed by self-supervised fine-tuning on specific downstream datasets. Moreover, text-based reports blankemeier2024merlin ; zhang2024generalist , which are commonly employed as auxiliary information in 3D CT pre-trained models, are often lacking for 3D MRI data, further limiting the development of comprehensive models. Therefore, the main challenge in building a general 3D MRI vision foundation model is collecting and curating a sufficiently large and diverse dataset. This dataset must cover a wide range of imaging modalities and hardware specifications to ensure the robustness and generalizability of the model.
[…]
2 Results
We present results from 25 downstream datasets across three types of evaluation tasks and two data modality settings. These downstream tasks are categorized into within-domain and out-of-domain tasks. Within-domain downstream tasks utilize the same data modalities and structures as those in the pre-training phase, including brain, breast, and prostate MRI. These tasks assess whether Triad has successfully learned structural and modality representations during pre-training, thereby improving performance on related tasks. Conversely, out-of-domain downstream tasks involve data modalities or organs different from those in the pre-training stage, such as liver CT or atrial MRI. These tasks evaluate whether the knowledge acquired by Triad during pre-training can be effectively transferred to and applied in new modalities or structures. Based on these two data modality settings, we evaluate three task types: 3D structure/tumor segmentation (Fig. 3 and Fig. 4), organ/cancer classification (Fig. 5), and 3D medical image registration (Fig. 6).
Figure 3: Study on within-domain 3D tumor segmentation. a. Image segmentation with encoder-decoder architecture by loading the weights of Triad. b. We compare the performance of Scratch, VoCo-SSL and Triad on 5 within-domain datasets based on 3 architectures: Swin-B/L/H. c. We select the nnUNet and Swin-Transformer-Base architectures, along with 3 different weight-loading strategies, and analyze their cross-effects on performance across 5 whitin-domain datasets. d. Consistent with the setting in subfig. c., the radar chart of each category shows the advantage of Triad in tumor segmentation.
2.1 3D organ/tumor segmentation
We first evaluated the effectiveness of Triad on 3D organ/tumor segmentation, a common tasks in medical image analysis. As shown in Fig. 3 (a), we initialize the encoder with the parameters learned during pre-training, while the decoder is randomly initialized. We evaluate Triad on 17 extensive 3D MRI and CT semantic segmentation datasets, including five MRI datasets for within-domain tasks: BraTS21(baid2021rsna, ), MSD(antonelli2022medical, )-BrainTumour, BreastDM(zhao2023breastdm, ), Prostate158(adams2022prostate158, ), and MSD-Prostate; and 12 datasets covering different organs or modalities for out-of-domain tasks: MM-WHS-MRI(zhuang2018multivariate, ), ATLAS-MRI(quinton2023tumour, ), Abdomen 1K(ma2021abdomenct, ), Kipa22(he2021meta, ), MSD-Pancreas, MSD-Liver, MSD-Heart, MSD-Hippocampus, MSD-Lung, MSD-HepaticVessel, MSD-Spleen, and MSD-Colon. We use the Dice Similarity Coefficient (DSC) as the primary evaluation metric, consistent with public benchmarks.
2.1.1 Influence of model parameter scale on model performance
It is widely believed that increasing model parameters enhances the performance of downstream tasks in foundation models (oquab2023dinov2, ; chen2024towards, ; ghesu2022contrastive, ; amadou2024echoapex, ). This trend has been observed in various domains, including natural images (oquab2023dinov2, ), X-rays (ghesu2022contrastive, ), and other medical imaging modalities(chen2024towards, ; amadou2024echoapex, ). However, some studies have reported contradictory findings, particularly in 3D CT imaging(wu2024voco, ) and vision language models (shi2025we, ; mei2024bigger, ). In this study, we quantitatively analyze the scaling behavior of Triad pre-training. We conduct a series of experiments with varying model architectures to systematically evaluate their impact on performance.
We specifically select the widely used SwinUNETR architecture (he2023swinunetr, ). The SwinUNETR encoder utilizes different variants of Swin Transformer (dosovitskiy2020image, ), including Swin-B (Base), Swin-L (Large), and Swin-H (Huge). We evaluate the impact of model parameter scaling on the within-domain 3D tumor segmentation task and compare it with Scratch and VoCo-SSL (wu2024voco, ). VoCo-SSL is a vision foundation model pre-trained on 160K 3D CT scans using a self-supervised model distillation scheme and is considered state-of-the-art in 3D medical imaging. Scratch denotes training from scratch without using any pre-trained weights.
As shown in Fig. 3 b (1) - b (5), the average DSC reported by Swin-B-Scratch across the five datasets is 77.76%. In comparison, Swin-B-VoCo-SSL achieves an average DSC of 79.31% (+1.55%), while Swin-B-Triad achieves 79.66% (+1.90%). These results indicate that pre-training the upstream encoder can markedly enhance the performance of downstream tasks. This finding is consistent with previous studies(wu2024voco, ; amadou2024echoapex, ). Notably, Swin-B-Triad outperforms Swin-B-VoCo-SSL by +0.35%, which can be attributed to Triad’s use of MRI data for pre-training, whereas VoCo-SSL is pre-trained on CT data. This suggests that greater alignment between the data modalities of upstream and downstream models leads to improved downstream performance. A key finding across 15 experiments comparing the Swin-B/L/H architectures is that 11 of these experiments indicate that increasing the size of model parameter does not consistently lead to performance improvements. This observation aligns with findings from VoCo-SSL(wu2024voco, ). A possible explanation is that excessive model parameters may lead to overfitting on small downstream datasets. Even with robust initial parameters from the upstream pre-trained model, downstream performance may still be adversely affected by overfitting.
2.1.2 Within-domain 3D tumor segmentation
The upstream 3D MRI data used for pre-training is derived from three organs: the brain, breast, and prostate. Consequently, the downstream within-domain 3D tumor segmentation task employs data from the same modality and organs to evaluate performance, assessing whether Triad has effectively learned structural and modality representations during pre-training. Specifically, we select two brain datasets: BraTS21 (baid2021rsna, ) and MSD-BrainTumour (antonelli2022medical, ), one breast dataset: BreastDM (zhao2023breastdm, ), and two prostate datasets: MSD-Prostate (antonelli2022medical, ) and Prostate158 (adams2022prostate158, ). We employ two widely used network architectures: nnUNet isensee2021nnu and SwinUNETR (Swin-B).
The nnUNet-Scratch achieves an average DSC of 79.68%, whereas nnUNet-Triad improved this to 81.13%, marking an increase of 1.45%. Considering the performance of the Swin-B encoder under the three initialization settings discussed earlier, it is evident that pre-trained parameters consistently outperform random initialization, regardless of the model structure (nnUNet or SwinUNETR) or the pre-training strategy (VoCo-SSL or Triad) employed. Furthermore, the results indicate that nnUNet generally outperforms SwinUNETR in segmentation tasks. A radar chart presenting DSC for each category in Fig. 3 (d). This chart demonstrates that Triad excels in segmenting fine-grained tumors. For example, the BraTS21 Tumor Core represents the core region of the tumor, which serves as the primary therapeutic target and excludes the edema area. The BraTS21 Enhancing Tumor represents the actively invasive tumor and serves as a key indicator of tumor grade and recurrence. Notably, Triad outperforms Swin-B-VoCo by +2.08% and +1.38% in both categories.
Figure 4: Study on out-of-domain organ/tumor segmentation. a. We select the nnUNet and Swin-Transformer-B architectures, along with three different weight loading strategies, and analyze their cross-effects on performance across six MSD CT datasets. b. Consistent with the setting of subfig. a., the radar chart shows the performance comparison of each category in MM-WHS-MRI, ATLAS-MRI, and MSD-Liver. c. Consistent with the setting of subfig. a., the radar chart shows the performance comparison of each category in Abdoman 1K, Kipa22, and MSD-Pancreas.
2.1.3 Out-of-domain organ/tumor segmentation
We further assess whether the knowledge acquired by Triad during pre-training can be effectively transferred to and applied in new modalities or anatomical organs. To achieve this, we select four MRI datasets from other organs: MSD-Heart, MSD-Liver, MM-WHS-MRI (zhuang2018multivariate, ), and ATLAS-MRI (quinton2023tumour, ). Additionally, we incorporated eight CT datasets: MSD-Hippocampus, MSD-Lung, MSD-Pancreas, MSD-HepaticVessel, MSD-Spleen, MSD-Colon, Abdomen 1K (ma2021abdomenct, ), and Kipa22 (he2021meta, ). We continue to use the nnUNet and SwinUNETR (Swin-B) architectures and employ three parameter initialization methods: training from scratch, VoCo-SSL, and Triad.
As shown in Fig. 4 (a), an interesting observation is that, when using the Swin-B architecture, the three initialization methods rank in performance as follows: VoCo (77.14%) > Scratch (74.34%) > Triad (73.09%)***We report the average DSC of experiments (1)-(6).. We also observed that VoCo-SSL was pre-trained on CT data containing more than 16 organs or tumors, covering the organs used in experiments (1)-(6). We believe this is the primary reason for its superior performance. Therefore, we conclude that pre-training maximizes performance when the data modality and organ of upstream and downstream tasks are consistent. Nevertheless, compared to both VoCo-SSL and training from scratch, nnUNet-Triad achieved an improvement of +2.02%, demonstrating that Triad can effectively generalize to other data modalities and organs. Fig. 4 (b) and (c) present radar charts for each class on both CT and MRI datasets. Notably, marked improvements are observed in tumor segmentation rather than organs. For example, on MSD-Liver Cancer, nnUNet-Triad outperforms nnUNet-Scratch by +14.77%, while Swin-B-Triad surpasses Swin-B-Scratch by +17.97%.
2.2 Organ/cancer classification
We next evaluate the performance of Triad on organ and cancer classification tasks. As shown in Fig. 5 (a), we initialize the encoder with the parameters learned during pre-training, apply an average pooling operation to the output of its final layer, and pass the resulting features through a two-layer linear classifier to predict the probability distribution of the categories. We evaluate Triad on five widely recognized 3D CT and MRI classification datasets, including two MRI datasets for within-domain classification: ADNI jack2008alzheimer and BreastDM zhao2023breastdm ; two CT datasets for out-of-domain classification: OrganMNIST3D yang2023medmnist and LUNA16 setio2017validation ; and one additional MRI dataset for out-of-domain classification: LLD-MMRI lou2024sdrformer . We use classification accuracy (Acc) as the primary evaluation metric.
2.2.1 Within-domain organ/cancer classification
As shown in Fig. 5 (c), we compare two architectures, 3D UNet and Swin-B, using four initialization methods: training from scratch, SwinUNETR cao2022swin , VoCo-SSL, and Triad. We observe that on both the ADNI and BreastDM datasets, Swin-B-Scratch achieves an average accuracy that is +4.25% higher than 3D UNet-Scratch. A similar trend is observed in the LLD-MMRI and OrganMNIST3D datasets. The only exception is the LUNA16 dataset, where 3D UNet-Scratch achieves an accuracy that is +0.73% higher than Swin-B-Scratch. These findings provide strong evidence that the Swin-B architecture is better suited for classification tasks.
Next, we compare the impact of three different pre-trained models on downstream performance. SwinUNETR is pre-trained on approximately 5K CT volumes, whereas VoCo-SSL utilizes 160K CT volumes. According to the reported accuracy, VoCo-SSL achieves an average accuracy that is +1.37% higher than SwinUNETR. Triad is pre-trained on 131K MR volumes and achieves an average accuracy that is +1.52% higher than VoCo-SSL. These results indicate that both the modality and scale of pre-training data positively impact downstream performance.
Figure 5: Study on organ/cancer classification. a. We use an encoder loaded with Triad weights and a two-layer linear classifier for classification tasks. b. Confusion matrices of the 5 datasets when using Swin-B-Triad as the encoder. The meaning of each category number is shown in Table 2. c. We select the 3D UNet and Swin-Transformer-Base architectures, along with 3 different weight loading strategies, and analyze their cross-effects on performance across 5 CT/MRI datasets. d. Consistent with the setting of subfig. c., we plot the ROC curve of each scheme on 4 datasets.
Figure 6: Study on 3D medical image registration. a. We adopt the TransMorph architecture, use Swin-Transformer Large as the encoder, and load the pre-trained weights of Scratch, Triad, and VoCo-SSL for regristation task. The bar chart on the right shows the average dice scores of the 3 weight loading methods on the 3 datasets. b. Under the setting of subfig. a., the visualization results of various registration methods in the IXI dataset. c. We adopt the SwinUNETR architecture, use Swin-Transformer Base as the encoder. The bar chart on the right shows the average dice scores of the 5 weight loading methods on the 2 datasets. d. Under the setting of subfig. c., the visualization results of various registration methods in the OASIS dataset. e. Under the setting of subfig. a., boxplots with Dice scores of various registration methods in the abdomen IXI dataset.
2.2.2 Out-of-domain organ/cancer classification
As illustrated in Fig. 5 (c), Triad achieves the highest performance in the organ classification task and ranks second in both lung nodule and liver lesion classification tasks. Notably, Triad still outperforms training from scratch by +1.02%, demonstrating its effectiveness in generalizing across diverse imaging modalities and organ types. Furthermore, we provide the confusion matrix for Swin-B-Triad across the five datasets. Fig. 5 b (1) shows that when Swin-B-Triad is applied to an out-of-domain classification task with a highly imbalanced category distribution, the model struggles to classify minority classes accurately. In the OrganMNIST3D classification task, Triad fails to distinguish categories 1–4 accurately. These findings suggest that while pre-trained parameters enhance overall downstream performance, addressing challenges such as data imbalance and hard example mining may require specialized sampling strategies or model architectures. Additionally, we present the ROC curves for four datasets in Fig. 5 (d). The ROC curves of all pre-trained models exhibit significant overlap, whereas models trained from scratch show markedly inferior performance, particularly on OrganMNIST3D and ADNI.
2.3 3D medical image registration
Finally, we evaluate the performance of Triad on the 3D medical image registration task. As illustrated in Fig. 6 (a) and (c), we explore two different parameter initialization strategies. In Fig. 6 (a), we employ the TransMorph (chen2022transmorph, ) architecture with a Swin-Transformer-L encoder, initializing it with pre-trained weights from Triad and VoCo-SSL. In Fig. 6 (c), we use the Swin-UNETR (hatamizadeh2021swin, ) architecture with a Swin-Transformer-B encoder, initializing it with pre-trained weights from Triad, VoCo-SSL, SuPreM, and SwinUNETR. The decoder remains unchanged from the original method and is randomly initialized. We evaluate Triad on three widely recognized 3D MRI registration datasets, including two brain datasets for within-domain registration: IXI Brain-Development_IXI_2019 and OASIS krentzel2023clem , as well as one cardiac dataset for out-of-domain registration: ACDC bernard2018deep . We use the Dice Similarity Coefficient (DSC) as the primary evaluation metric and report the best results after fine-tuning for 200 epochs.
2.3.1 Comparison of pre-training strategies in TransMorph and SwinUNETR for 3D medical image registration
Fig. 6 (a) illustrates a bar chart depicting the DSCs for each dataset. TransMorph-Scratch achieves average DSCs of 73.76%, 86.79%, and 74.81% on IXI, OASIS, and ACDC, respectively. When employed Triad pre-trained weights, the DSCs are 73.91% (+0.15%) on IXI and 86.92% (+0.13%) on OASIS, but decrease to 74.57% (-0.24%) on ACDC. Similarly, VoCo-SSL pre-training results in DSCs of 73.44% (-0.32%), 86.98% (+0.19%), and 74.72% (-0.09%) on IXI, OASIS, and ACDC, respectively. The performance of TransMorph under these three initialization strategies indicates that pre-trained parameters do not consistently yield improvements over random initialization in 3D medical image registration. This observation applies to both within-domain (IXI, OASIS) and out-of-domain (ACDC) tasks, with some cases even exhibiting marginal performance declines.
Fig. 6 (c) presents DSC performance under the SwinUNETR architecture, comparing five initialization strategies: Scratch, SwinUNETR, SuPreM, VoCo-SSL, and Triad. Specifically, Swin-B-Scratch achieves DSCs of 72.60% on IXI and 81.80% on OASIS. Swin-B-VoCo achieves DSC scores of 73.60% (+1.00%) on IXI and 84.40% (+2.60%) on OASIS. Notably, Triad proves to be the most effective pre-training method, achieving DSC scores of 73.70% (+1.10%) on IXI and 88.70% (+6.90%) on OASIS. These improvements are particularly pronounced on the OASIS dataset, where Triad outperforms other initialization methods by a significant margin.
By integrating these findings with the TransMorph results in Fig. 6 (a), we observe that partially loading encoder weights, as done in TransMorph, while randomly initializing the remaining parameters, may introduce inconsistencies, potentially limiting the benefits of pre-training. In contrast, SwinUNETR employs a fully pre-trained encoder, thereby eliminating random initialization in that module. This allows the network to leverage pre-trained features more effectively, leading to substantial improvements in 3D medical image registration.
2.3.2 Impact of initialization method on registration performance
Fig. 6 (b) presents the visualization results for the IXI dataset using the TransMorph architecture. Regardless of the initialization method used for fine-tuning, the observed improvements in the mask appear similar, with no substantial enhancements detected. In contrast, Fig. 6 (d) presents the visualization of the OASIS dataset using the SwinUNETR architecture, where improvements in the mask (indicated by the red arrow) are noticeable. These visualizations provide intuitive evidence supporting our previous assertion that incomplete pretraining initialization of the encoder may lead to model confusion.
Furthermore, Fig. 6 (e) depicts the DSC distributions for the IXI dataset. SwinUNETR initialized with Triad weights achieves the highest registration performance across most organs, including the Thalamus, Cerebral White Matter, Cerebellar White Matter, Pallidum, Caudate, Lateral Ventricle, Hippocampus, Third Ventricle, Fourth Ventricle, and Amygdala. Notably, this superior performance can be attributed to the inclusion of abundant 3D MRI brain organ data in the upstream pretraining tasks. Regardless of the initialization weights, the DSC and registration performance for the choroid plexus remains low, likely due to its complex anatomical attachments to surrounding structures and its diffuse morphological characteristics.
3 Discussion
In this study, we constructed a large-scale 3D MRI dataset, known as TriadMR-131K, which consists of 131,170 volumes from 19,721 patients across 36 clinical datasets. This extensive collection includes a diverse collection of 3D MRI data from three organs, including the brain, breast, and prostate. It includes modalities such as T1-w, T2-w, FLAIR, DWI-MRI, fMRI, and DCE-MRI. Additionally, we extract imaging descriptions from the metadata of each 3D volume. These descriptions detail the imaging modality and associated device parameters. Using this dataset, we develop Triad, a vision foundation model tailored for 3D MRI. Triad employs widely used autoencoder architecture to learn robust representations and incorporates organ-independent imaging descriptions to constrain the semantic distribution of the visual modality. We evaluate Triad on three tasks, including organ and tumor segmentation, organ and cancer classification, and medical image registration. These tasks are assessed across two data modalities—within-domain and out-of-domain— across 25 downstream datasets. By initializing models with Triad’s pre-trained weights, nnUNet-Triad improves segmentation performance by 6.88% over nnUNet-Scratch across 17 datasets. Swin-B-Triad achieves a 3.97% improvement over Swin-B-Scratch in classification tasks across five datasets. SwinUNETR-Triad improves by 4.00% compared to SwinUNETR-Scratch in registration tasks across three datasets. Triad outperforms baseline models across all downstream tasks and exceeds existing state-of-the-art models in most cases. Overall, Triad’s seamless adaptability across downstream tasks highlights its potential as a versatile and efficient tool for diverse clinical applications. This adaptability paves the way for enhanced diagnostic accuracy in 3D MRI.
Our findings highlight several key observations regarding the effectiveness of pretraining:
  • Most vision foundation models for medical image analysis have been pre-trained on large-scale 3D CT datasets (tang2022self, ; ye2024continual, ; zhuang2024mim, ; jiang2023anatomical, ), highlighting the prevalence of CT-based self-supervised learning in this domain. However, our experimental results indicate that pre-training on modality-specific upstream data is more effective. This is demonstrated by the superior performance of Triad, which was pre-trained on MRI data, in 3D MRI-based segmentation, classification, and registration tasks. This finding underscores a critical gap in current research, namely the absence of large-scale 3D MRI pre-training datasets for foundation models. By assembling a diverse collection of 3D MRI data, our work addresses this gap and establishes a more appropriate pre-training paradigm for 3D MRI-based medical imaging tasks. Furthermore, although this study employs an autoencoder (AE) architecture that is not the most advanced in self-supervised learning, our comparison with VoCo-SSL, a state-of-the-art distilled model, suggests that even a relatively simple AE-based framework can outperform more complex models under certain conditions. This highlights an important insight: data quality and task-specific alignment are more critical than model complexity, suggesting that the future of medical AI should emphasize comprehensive and representative pre-training datasets rather than focusing solely on architectural advancements.
  • The experimental results of Triad pre-training highlight the dual impact of model parameter scale and task alignment. In the context of 3D organ and tumor segmentation, pre-trained encoders consistently outperform their randomly initialized counterparts across various architectures. However, increasing the scale of the model parameters does not consistently lead to improved performance, as evidenced by 11 out of 15 experiments showing no marked gain. This suggests that while larger models theoretically have greater representational capacity, the limited size of downstream medical imaging datasets may lead to overfitting, offsetting the benefits of pre-training. Furthermore, Triad, which is pre-trained on MRI data, surpasses the CT-based VoCo-SSL model in MRI-related tasks, underscoring the critical role of modality alignment in transfer learning for medical imaging. Notably, Triad demonstrates superior segmentation performance in fine-grained tumor regions, such as the enhancing tumor component in BraTS21. This suggests that modality alignment not only affects overall segmentation accuracy but also improves the precise delineation of clinically relevant tumor subregions.
  • In out-of-domain tasks, Triad exhibits strong performance in MRI-based applications but lags behind VoCo-SSL in CT-based segmentation. Although Triad does not surpass VoCo-SSL in CT tasks, it significantly outperforms models trained from scratch and achieves notable improvements in certain tumor segmentation tasks, such as liver cancer. These results suggest that the benefits of pre-training are more pronounced for lesion recognition than for anatomical organ segmentation. Additionally, in classification tasks, Swin-B consistently outperforms 3D UNet, suggesting that Transformer-based architectures may be more suitable for medical image classification. However, in highly imbalanced datasets, Triad-pretrained models struggle with minority class discrimination, underscoring the limitations of pre-training alone in mitigating data imbalance. These findings suggest that future work should incorporate advanced sampling strategies or hard example mining techniques to further improve model generalization across diverse medical imaging tasks.
  • In registration tasks using TransMorph, loading pre-trained weights from Triad and VoCo-SSL resulted in only marginal improvements in within-domain registration tasks, such as IXI and OASIS. However, in some cases, including the out-of-domain ACDC dataset, performance slightly declined compared to random initialization. These results suggest that partially loading encoder weights, as in TransMorph, may introduce inconsistencies that impede optimal feature learning during fine-tuning. In contrast, SwinUNETR, which employs a fully pre-trained encoder, achieved more substantial improvements when initialized with Triad weights. This effect was particularly evident in the OASIS dataset, where the DSC increased markedly by 6.90% compared to random initialization. These findings indicate that leveraging a fully pre-trained encoder enables more effective feature transfer, thereby improving registration accuracy. Furthermore, the superior performance of Triad over other pretraining approaches underscores the advantages of task-specific pretraining strategies tailored for 3D MRI data.
Despite the strengths of our study, there are several limitations in both our model and methodology. First, the text data used for pre-training in Triad-131K primarily focuses on device parameters and imaging modalities rather than image content. This restricts the extension of the single-modality vision foundation model into a vision-language foundation model. Second, the computational demands of pre-training such a large-scale 3D MRI dataset prevent us from employing ensemble methods and fixed-step validation, both of which could further enhance the performance of downstream tasks. We retain model parameters only at iteration 20,000. Third, fine-tuning across a broad range of downstream tasks is highly resource-intensive in terms of manpower, computational cost, and time. Following VoCo’s approach, we report results solely on fold 0 instead of conducting multi-fold cross-validation, which may introduce variability due to data distribution effects. Fourth, additional exploration is required to optimize Triad and enhance its performance on downstream tasks. For instance, in the registration task, we initialize TransMorph with Triad’s pre-trained weights and subsequently fine-tune the model. The training process involves balancing multiple loss functions; however, we use default parameters without exploring a broader optimization space. Future research should explore more effective parameter configurations and optimization strategies.
Future research can focus on the following aspects. First, improving data quality. Despite having implemented various preprocessing pipelines, the presence of low-quality cases remains a challenge. Numerous studies have emphasized the crucial role of data quality in pre-training (yang2024freemask, ; kirillov2023segment, ). In future work, we plan to design robust automated screening pipelines to screen each volume and enhance data quality. Second, assigning structured electronic health reports to each 3D MRI volume in the Triad-131K dataset to facilitate pre-training for vision-language foundation models. Additionally, expanding downstream tasks to include assistive report generation, visual question answering, and cross-modal medical image retrieval. Lastly, the current pre-training strategy is not limited to MRI scans. Moving forward, we plan to integrate Triad-131K with the largest available CT, PET, X-ray, and ultrasound datasets to pre-train foundation models that generalize effectively across a broad range of radiology tasks, rather than being confined to specific imaging modalities.
4 Methods
The following sections are structured are organized to provide a comprehensive overview of our methodologies and findings. Initially, we will introduce the data curation process for TriadMR-131K, as well as the protocol established for Triad pre-training. Following this, we will articulate the implementation strategies employed in three distinct experimental paradigms: 3D organ and tumor Segmentation, organ and cancer classification, and 3D medical image registration.
4.1 Pretraining
Curation of TriadMR-131K dataset. To ensure the quality and diversity of data for model pretraining, we curated a large-scale 3D MRI dataset of 131,170 3D volumes derived from 19,721 patients across 36 clinical datasets. TriadMR-131K comprises a diverse collection of 3D MRI data spanning three organs (brain, breast, and prostate), featuring modalities such as T1-w, T2-w, FLAIR, DWI-MRI, fMRI, DCE-MRI. To standardize all sub-datasets, we used the same preprocessing protocol for all organs: we used the dicom2niftihttps://github.com/icometrix/dicom2nifti package to convert all DICOM-format 2D slice collections into NIfTI-format 3D volumes. For 4D volume data, such as DCE-MRI, we took the $\frac{(t-1)}{2}$th or $\frac{(t+1)}{2}$th 3D slice to replace the original 4D data, where $t$ denotes the $t$th time step. All corrupted volumes were deleted during the conversion process. Next, we reformatted all MRI scans so that the first axis points from left to right, the second from posterior to anterior, and the third from inferior to superior. We then resampled the images to a 1 mm resolution using bilinear interpolation. We also resized all images to (256,256,128) using trilinear interpolation. To save storage space, we stored most of the 3D volume data types as UINT16 and the rest as Float32. Finally, the brain MRI data involved 51,112 series from 37,436 examinations of 12,994 patients; the breast MRI data involved 46,116 series from 8,180 examinations of 3,834 patients; the prostate MRI data involved 33,942 series from 9,941 examinations of 4,639 patients. The statistical information of the 36 datasets is shown in Table 1. Note that due to deletions during the conversion process, the numbers in the table are usually lower than the officially published figures. In addition, we extracted the imaging description from the metadata of each 3D volume, which describes the imaging modality and related device parameters. Since it is organ-independent information, it helps to adjust the positional relationship of each modality in the semantic space, thereby improving the generalization ability of the model. We tried to avoid any overlap between the datasets used in pre-training and all downstream evaluation sets to minimize the risk of data contamination.
Protocol of Triad pre-training. Triad uses nnUNet (31M)isensee2021nnu and SwinTransformerhatamizadeh2021swin for the image encoder. Furthermore, SwinTransformer is expanded into Swin-B (72.8M), Swin-L (290.4M), and Swin-H (11.6B) according to feature sizes of 48, 96, and 192 to study the parameter scaling law (Fig. 2b). We use GTR-T5-Large ni2021large as the text encoder instead of CLIP radford2021learning because the text embedding of T5 can provide a more semantically nuanced distribution and is suitable as a supervisory signal to guide the alignment of visual modality distributions to the semantic space, rather than just as a representation of relative similarity in contrastive learning. This means that the parameters of the text encoder are frozen and do not participate in gradient updates. In the pretraining stage, we use an upsampling 3D CNN as the image decoder for the self-supervised reconstruction task. In the downstream task, we only keep the parameters of the image encoder and replace the image decoder with the task adapter. Triad’s pretraining uses L1 loss as the reconstruction loss and log-ratio loss kim2019deep to align the distribution of the visual modality with that of the textual modality. To prevent the log-ratio loss from dominating the optimization process, its weight is set to 0.01. The architectural hyperparameters of the models involved are shown in Table 3.
4.2 3D organ/tumor segmentation
4.2.1 Curation of segmentation datasets
Within-domain segmentation datasets. The BraTS21 dataset (baid2021rsna, ), released as part of the 2021 RSNA-ASNR-MICCAI Brain Tumor Segmentation Challenge, consists of multi-institutional, multi-modal MRI scans (T1-w, T1 postcontrast, T2-w, FLAIR) from patients with glioblastoma or lower-grade glioma. Each case includes expert annotations delineating tumor subregions (enhancing tumor, edema, and necrotic core). The core challenge portion provides 1,251 labeled scans for training. The MSD–BrainTumour dataset (antonelli2022medical, ) (Task 01 in the Medical Segmentation Decathlon) includes 484 preoperative multi-modal MRI scans (T1-w, T1 postcontrast, T2-w, FLAIR) sourced primarily from earlier BraTS collections. For BreastDM (zhao2023breastdm, ), the original publication reports a new breast MRI dataset (with dynamic contrast-enhanced volumes) consisting of 232 scans. The Prostate158 dataset (adams2022prostate158, ) offers 158 MRI scans (T2, ADC, DFI) with detailed prostate annotations. Finally, the MSD–Prostate dataset (antonelli2022medical, ) (Task 05 in the MSD) contains 32 T2-w, ADC map, and DFI scans with corresponding prostate zonal annotations (central gland and peripheral zone).
Out-of-domain segmentation datasets. The MM-WHS–MRI subset (zhuang2018multivariate, ) from the Multimodality Whole Heart Segmentation challenge consists of around 20 annotated 3D MRI volumes of the heart. ATLAS–MRI (quinton2023tumour, ) is a publicly available dataset of contrast-enhanced MRI for hepatocellular carcinoma (HCC), which consists of 60 scans. The Abdomen-1K (ma2021abdomenct, ), released as part of the MICCAI FLARE 2022 Playground subtask 1, includes a training set adapted from MSD Pancreas (281 cases) and NIH Pancreas (80 cases), where all 361 CT scans are from the portal phase. Kipa22 (he2021meta, ) comes from the Kidney PArsing Challenge 2022, and its goal is to segment 3D kidneys, renal tumors, arteries, and veins. It released 70 training sets with detailed annotations. Lastly, the remaining MSD tasks, namely MSD-Pancreas (281 training scans), MSD-Liver (131 training scans), MSD-Heart (20 training scans), MSD-Hippocampus (260 training scans), MSD-Lung (63 training scans), MSD-HepaticVessel (303 training scans), MSD-Spleen (41 training scans), and MSD-Colon (126 training scans), do not provide official validation sets. For all the above datasets, we keep the same split method as provided by VoCo (wu2024voco, ).
4.2.2 Fully supervised finetuning with nnUNet framework
In our image segmentation experiments, we adopt nnUNetv2 as a unified framework to ensure consistent data preprocessing for fair comparisons across different models. Within this framework, we have implemented Swin-Transformer Base/Large/Huge networks, thereby aligning the training protocols. For each publicly available dataset with detailed annotations, nnUNetv2’s built-in code is used to perform a 5-fold cross-validation split. Note that, in order to make a fair comparison with VoCo, we report the 0th fold in the 5-fold cross-validation. Throughout training, nnUNet models are trained for 300 epochs, while Swin-Transformer models are trained for 150 epochs, and we select the model with the highest validation performance for final evaluation. We employ an SGD optimizer with an initial learning rate of 0.01, following nnUNet’s default decay schedule. VoCo-SSL pretrained weights are sourced from the code libraryhttps://github.com/Luffy03/Large-Scale-Medical. Because our experimental setup closely matches that of VoCo, some of the results reported here are derived from the extended version of the original publication (wu2024voco, ).
4.3 Organ/cancer classification
4.3.1 Curation of classification datasets
Within-domain classification datasets. The ADNI dataset (jack2008alzheimer, ) (Alzheimer’s Disease Neuroimaging Initiative) is a longitudinal, multi-center, observational study that includes thousands of participants, from cognitively normal (CN) to those with mild cognitive impairment (MCI) or Alzheimer’s disease (AD). In this study, we use a dataset consisting of participants who have screening, 6-month, 1-year, 18-month (MCI only), 2-year, and 3-year (normal and MCI only) scans, which is called “ADNI1_Complete 3Yr 1.5T,” totaling 2,182 samples. Consistent with the literature majee2024enhancing , the training set, validation set, and test set contain 1,526, 326, and 330 samples, respectively. We use NPPY he2023neural and its available pre-trained weights to convert raw MRI scans into uniformly sized skull-stripped, intensity-normalized brain volumes in standard coordinate space, and then reshape them into a smaller dimension of 96×96×96. The BreastDM dataset (zhao2023breastdm, ) provides dynamic contrast-enhanced (DCE) breast MRI scans for lesion classification, containing 85 benign samples and 147 malignant samples. We adopt the same dataset split scheme as in the segmentation task.
Out-of-domain classification datasets. The OrganMNIST3D dataset (yang2023medmnist, ) is part of MedMNIST v2 and contains more than 1,700 3D CT volumes of 11 organs for classification. Its official distribution includes dedicated training sets (972 volumes), validation sets (161 volumes), and test sets (610 volumes). The samples in OrganMNIST3D are available in 28×28×28 and 64×64×64 versions, and we use the latter for evaluation. The LUNA16 dataset (setio2017validation, ), derived from the LIDC/IDRI collection, contains 888 thoracic CT scans for lung nodule analysis. LUNA16 includes a total of 551,065 candidate nodules, of which 1,120 nodules are detected as positive, represented by 1, and the rest are represented by 0. The full dataset is divided into 10 subsets, and we use subsets 0-5 as training sets, subset 6 as the validation set, and subsets 7-9 as test sets. Finally, the LLD-MMRI dataset (lou2024sdrformer, ) contains 498 annotated multi-stage liver lesions from the same number of patients. The lesions are classified into seven categories: hepatocellular carcinoma (HCC), intrahepatic cholangiocarcinoma (ICC), hepatic metastasis (HM), hepatic cyst (HC), hepatic hemangioma (HH), focal nodular hyperplasia (FNH), and hepatic abscess (HA). The dataset has been pre-partitioned into a training set (316 lesions), a validation set (78 lesions), and a test set (104 lesions).
4.3.2 Fully supervised finetuning with linear classifier
As shown in Fig. 5 (a), we use the parameters saved in the pretraining phase as the initial parameters of the encoder, perform an average pooling operation on the output of the last layer of the encoder, and then input it into a two-layer linear classifier to predict the probability distribution of the category. In classification experiments, we set the ADNI dataset input size to 96×96×96, while all other datasets are resized to 64×64×64. Classifiers based on the Swin-Transformer are trained for 150 epochs, and those based on the 3D UNet are trained for 300 epochs. In each experiment, we report the best result. We employ a learning rate of 1e-3, with the Adam optimizer, following a cosine decay schedule. Additionally, the first 5 epochs are used for warmup to stabilize training.
4.4 3D medical image registration
4.4.1 Curation of registration datasets
Within-domain registration datasets. The IXI dataset (Brain-Development_IXI_2019, ) consists of over 576 T1-weighted brain MRI scans from healthy volunteers collected at three different hospitals in London. Following the TransMorph (chen2022transmorph, ) protocol, we use 403 scans as the training set, 58 as the validation set, and 115 as the test set. The volumes are cropped to 160×192×224. Thirty annotated structures were used for evaluation. The OASIS dataset (krentzel2023clem, ) (Open Access Series of Imaging Studies) includes 413 T1-weighted brain MRI scans from participants aged 18 to 96, with both healthy controls and patients exhibiting mild to moderate cognitive impairment. The original MR volumes are preprocessed using FreeSurfer fischl2012freesurfer , which includes spatial normalization, skull stripping, affine transformations, and automatic structural segmentation. Following the TransMorph (chen2022transmorph, ) protocol, we use 394 scans as the training set and 19 scans as the validation set. Since there is no test set available, we employed the validation set for evaluation. The volumes are cropped to 160×192×224. 35 structures are used as ground truths to evaluate the performance.
Out-of-domain registration datasets. The ACDC dataset (bernard2018deep, ) (Automated Cardiac Diagnosis Challenge) comprises 150 cardiac MRI scans in short-axis view, covering subjects with various heart conditions. The original challenge reserves 100 scans for training and 50 for testing. The volumes are cropped to 160×192×224.
4.4.2 Fine-tuning TransMorph/SwinUNETR for image registration
For fine-tuning of TransMorph, we replaced the “Transformer Encoder” in the original framework with our own Swin-Transformer Encoder, loading the weights of Triad-L. The rest of the components, such as “CNN Decoder,” “Affine Network,” and “Spatial Transform,” are randomly initialized. For fine-tuning of SwinUNETR, we load the weights of Triad-B into the encoder and randomly initialize the UNETR decoder. The pre-trained weights of SwinUNETR and SupreM are obtained from the code repository provided by VoCo. Due to limited resources, we only fine-tune for 200 epochs on each set of experiments and select the best-performing results for reporting. The Adam optimizer is used for fine-tuning, and the batch size was 1. The learning rates for OASIS and IXI are 0.00005, while the learning rate for ACDC is 0.0001. The remaining parameters, such as the type of loss function and weight factor, remain consistent with the default settings in the code provided by TransMorph.
4.5 Computing hardware and software
We use pydicom 3.0.1 and dicom2nifti 2.5.0 for 2D slice sequences and 3D volume data preprocessing. We use Python 3.10.13 for all experiments and analyses in the study. For the pretraining stage, we use the AdamW loshchilov2017decoupled optimizer with an initial learning rate of 1e-6, coupled with a cosine learning rate scheduler. The learning rate decays to zero over 200,000 steps, with a warm-up phase during the first 1,000 steps. We use two 80-GB NVIDIA A100 GPUs configured for multi-GPU training using DistributedDataParallel (DDP) as implemented by the framework PyTorch (version 2.5.1, CUDA 12.4), with a batch size of 8. We do not divide the data for TriadMR-131K but use all the data to pretrain Triad and then save the model at 200,000 steps to serve as the initial parameters for downstream tasks. For fine-tuning and validation of downstream tasks, we use the repository provided by VoCo v2 wu2024voco ; wu2024large (https://github.com/Luffy03/Large-Scale-Medical) and TransMorph chen2022transmorph (https://github.com/junyuchen245/TransMorph_Transformer_for_Medical_Image_Registration), respectively.
4.6 Evaluation metrics
We used several evaluation metrics to thoroughly assess the capabilities of our Triad model across different tasks. Accuracy is a primary metric used for evaluating the performance in medical-image classification, it is defined as the ratio of the number of correctly predicted samples to the total number of samples:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
where TP (True Positives) denotes the number of samples correctly predicted as positive by the model; TN (True Negatives) denotes the number of samples correctly predicted as negative by the model; FP (False Positives) denotes the number of negative samples incorrectly predicted as positive by the model; FN (False Negatives) denotes the number of positive samples incorrectly predicted as negative by the model.
Dice Similarity Coefficient (DSC) is used to measure the overlap between two sets, which is widely used in medical image segmentation tasks:
$\text{DSC}=\frac{2|X\cap Y|}{|X|+|Y|}$ (2)
where $X$ is the pixel set of the predicted segmentation result; $Y$ is the pixel set of the ground truth segmentation; $|X\cap Y|$ denotes the number of pixels contained in the intersection of $X$ and $Y$; $|X|$ and $|Y|$ denote the number of pixels of $X$ and $Y$ respectively. Equivalently, the DSC can be calculated based on the pixel category, which is expressed in pixel-by-pixel binary form by the predicted label $P$ and the ground truth label $G$:
$\text{DSC}=\frac{2\sum_{i}P_{i}G_{i}}{\sum_{i}P_{i}+\sum_{i}G_{i}}$ (3)
where $P_{i}$ is the predicted label for the $i$-th pixel value; $G_{i}$ is the $i$-th pixel value of the ground truth label. The value of DSC ranges between 0 and 1, where 1 indicates perfect overlap and 0 indicates no overlap.
5 Data Availability
All data in this study are publicly available and can be accessed from: fastMRI-Brain knoll2020fastmri ; zbontar2018fastmri (https://fastmri.med.nyu.edu/), UCSF-PDGM calabrese2022university (https://www.cancerimagingarchive.net/collection/ucsf-pdgm/), ACRIN-DSC-MR-Brain kinahan2019data (https://www.cancerimagingarchive.net/collection/acrin-dsc-mr-brain/), TCGA-GBM scarpace2016cancer (https://www.cancerimagingarchive.net/collection/tcga-gbm/), UPENN-GBM bakas2021multi (https://www.cancerimagingarchive.net/collection/upenn-gbm/), IvyGAP pati2020multi (https://www.cancerimagingarchive.net/analysis-result/ivygap-radiomics/), ACRIN-FMISO-Brain kinahan2018data (https://www.cancerimagingarchive.net/collection/acrin-fmiso-brain/), IXI (https://brain-development.org/ixi-dataset/), CPTAC-GBM CPTACGBM2018 (https://www.cancerimagingarchive.net/collection/cptac-gbm/), ReMIND juvekar2023brain (https://www.cancerimagingarchive.net/collection/remind/), REMBRANDT scarpace2019data (https://www.cancerimagingarchive.net/collection/rembrandt/), VESTIBULAR-sHWANNOMA Kujawa2023_Vestibular_Schwannoma (https://www.cancerimagingarchive.net/collection/vestibular-schwannoma-mc-rc/), GLIS-RT shusharina2021glioma (https://www.cancerimagingarchive.net/collection/glis-rt/), Meningioma-SEG-CLASS vassantachart2023segmentation (https://www.cancerimagingarchive.net/collection/meningioma-seg-class/), ISPY2 newitt2021acrin ; Li2022_ISPY2 (https://www.cancerimagingarchive.net/collection/ispy2/), ACRIN-6698 newitt2021acrin (https://www.cancerimagingarchive.net/collection/acrin-6698/), ACRIN-Contralateral-Breast-MR Kinahan2021 (https://www.cancerimagingarchive.net/collection/acrin-contralateral-breast-mr/), EA1141 comstock2023abbreviated (https://www.cancerimagingarchive.net/collection/ea1141/), Advanced-MRI-Breast-Lesions Daniels2024 (https://www.cancerimagingarchive.net/collection/advanced-mri-breast-lesions/), ISPY1 newitt2016multi (https://www.cancerimagingarchive.net/collection/ispy1/), TCGA-BRCA lingle2016cancer (https://www.cancerimagingarchive.net/collection/tcga-brca/), Breast-MRI-NACT-Pilot newitt2016single (https://www.cancerimagingarchive.net/collection/breast-mri-nact-pilot/), QIN Breast DCE-MRI huang2014variations (https://www.cancerimagingarchive.net/collection/qin-breast-dce-mri/), fastMRI-Breast knoll2020fastmri ; zbontar2018fastmri (https://fastmri.med.nyu.edu/), BREAST-DIAGNOSIS bloch2015breast (https://www.cancerimagingarchive.net/collection/breast-diagnosis/), PROSTATEx litjens2017prostatex (https://www.cancerimagingarchive.net/collection/prostatex/), BIMCV-Prostate Alzate-Grisales2024 (https://bimcv.cipf.es/bimcv-projects/prostate/), PI-CAI Challenge saha2024artificial (https://pi-cai.grand-challenge.org/), Prostate-MRI-US-Biopsy natarajan2020prostate (https://www.cancerimagingarchive.net/collection/prostate-mri-us-biopsy/), CPTAC-UCEC CPTAC2019 (https://www.cancerimagingarchive.net/collection/cptac-ucec/), fastMRI-Prostate knoll2020fastmri ; zbontar2018fastmri (https://fastmri.med.nyu.edu/), PROSTATE-DIAGNOSIS bloch2015data (https://www.cancerimagingarchive.net/collection/prostate-diagnosis/), Prostate Fused-MRI-Pathology madabhushi2016fused (https://www.cancerimagingarchive.net/collection/prostate-fused-mri-pathology/), PROSTATE-MRI choyke2016data (https://www.cancerimagingarchive.net/collection/prostate-mri/), TCGA-PRAD Zuley2016 (https://www.cancerimagingarchive.net/collection/tcga-prad/), Prostate-3T litjens2015data (https://www.cancerimagingarchive.net/collection/prostate-3t/), MSD Challenge antonelli2022medical (https://decathlon-10.grand-challenge.org/), BraTs21 baid2021rsna (http://braintumorsegmentation.org/), BreastDM zhao2023breastdm (https://github.com/smallboy-code/Breast-cancer-dataset), Prostate158 adams2022prostate158 (https://zenodo.org/records/6481141), MM-WHS-MRI zhuang2018multivariate (https://zmiclab.github.io/zxh/0/mmwhs/), ATLAS-MRI quinton2023tumour (https://atlas-challenge.u-bourgogne.fr/), Abdoman 1K ma2021abdomenct (https://github.com/JunMa11/AbdomenCT-1K?tab=readme-ov-file), Kipa22 he2021meta (https://kipa22.grand-challenge.org/), OASIS marcus2007open (https://sites.wustl.edu/oasisbrains/), ACDC bernard2018deep (https://www.creatis.insa-lyon.fr/Challenge/acdc/databases.html), OrganMNIST3D yang2023medmnist (https://github.com/MedMNIST/MedMNIST/tree/main), LUNA16 setio2017validation (https://luna16.grand-challenge.org/Data/), LLD-MMRI lou2024sdrformer (https://github.com/LMMMEng/LLD-MMRI-Dataset), ANDI jack2008alzheimer (https://adni.loni.usc.edu/data-samples/adni-data/neuroimaging/mri/mri-image-data-sets/).
6 Code Availability
The pretrained and fine-tuned models, as well as source code for training, inference and data preprocessing, can be accessed at https://github.com/wangshansong1/Triad.
Supplementary Materials: Tables and Figures
Table 1: Datasets used in Triad for pre-training with details.
Organ Dataset Series Number Subjects Studies Availability
Brain fastMRI-Brain 23,082 8,165 23,082 https://fastmri.med.nyu.edu/
UCSF-PDGM 6,998 501 6,998 https://www.cancerimagingarchive.net/collection/upenn-gbm/
ACRIN-DSC-MR-Brain 5,093 123 547 https://www.cancerimagingarchive.net/collection/acrin-dsc-mr-brain/
TCGA-GBM 3,547 255 520 https://www.cancerimagingarchive.net/collection/tcga-gbm/
UPENN-GBM 3,129 627 3,109 https://www.cancerimagingarchive.net/collection/upenn-gbm/
IvyGAP 2,769 39 385 https://www.cancerimagingarchive.net/analysis-result/ivygap-radiomics/
ACRIN-FMISO-Brain 2,071 42 187 https://www.cancerimagingarchive.net/collection/acrin-fmiso-brain/
IXI 1,090 581 1,090 https://brain-development.org/ixi-dataset/
CPTAC-GBM 958 63 127 https://www.cancerimagingarchive.net/collection/cptac-gbm/
ReMIND 638 114 227 https://www.cancerimagingarchive.net/collection/remind/
REMBRANDT 499 90 97 https://www.cancerimagingarchive.net/collection/rembrandt/
VESTIBULAR-sHWANNOMA 484 242 484 https://www.cancerimagingarchive.net/collection/vestibular-schwannoma-mc-rc/
GLIS-RT 431 230 431 https://www.cancerimagingarchive.net/collection/glis-rt/
Meningioma-SEG-CLASS 323 89 152 https://www.cancerimagingarchive.net/collection/meningioma-seg-class/
Breast ISPY2 18,494 714 2,648 https://www.cancerimagingarchive.net/collection/ispy2/
ACRIN-6698 8,668 385 1,148 https://www.cancerimagingarchive.net/collection/acrin-6698/
ACRIN-Contralateral-Breast-MR 5,906 788 875 https://www.cancerimagingarchive.net/collection/acrin-contralateral-breast-mr/
EA1141 3,526 500 953 https://www.cancerimagingarchive.net/collection/ea1141/
Advanced-MRI-Breast-Lesions 3,435 632 632 https://www.cancerimagingarchive.net/collection/advanced-mri-breast-lesions/
ISPY1 3,292 220 834 https://www.cancerimagingarchive.net/collection/ispy1/
TCGA-BRCA 867 137 156 https://www.cancerimagingarchive.net/collection/tcga-brca/
Breast-MRI-NACT-Pilot 694 64 201 https://www.cancerimagingarchive.net/collection/breast-mri-nact-pilot/
QIN Breast DCE-MRI 632 10 20 https://www.cancerimagingarchive.net/collection/qin-breast-dce-mri/
fastMRI-Breast 600 300 600 https://fastmri.med.nyu.edu/
BREAST-DIAGNOSIS 290 84 113 https://www.cancerimagingarchive.net/collection/breast-diagnosis/
Prostate PROSTATEx 17,313 346 351 https://www.cancerimagingarchive.net/collection/prostatex/
BIMCV-Prostate 6,397 1,501 1,531 https://bimcv.cipf.es/bimcv-projects/prostate/
PI-CAI Challenge 5,995 1,476 5,995 https://pi-cai.grand-challenge.org/
Prostate-MRI-US-Biopsy 1,994 837 1,184 https://www.cancerimagingarchive.net/collection/prostate-mri-us-biopsy/
CPTAC-UCEC 778 36 38 https://www.cancerimagingarchive.net/collection/cptac-ucec/
fastMRI-Prostate 625 312 625 https://fastmri.med.nyu.edu/
PROSTATE-DIAGNOSIS 261 89 89 https://www.cancerimagingarchive.net/collection/prostate-diagnosis/
Prostate Fused-MRI-Pathology 233 28 28 https://www.cancerimagingarchive.net/collection/prostate-fused-mri-pathology/
PROSTATE-MRI 158 26 26 https://www.cancerimagingarchive.net/collection/prostate-mri/
TCGA-PRAD 124 10 10 https://www.cancerimagingarchive.net/collection/tcga-prad/
Prostate-3T 64 64 64 https://www.cancerimagingarchive.net/collection/prostate-3t/
Total TriadMR-131K 131,170 19,721 55,557
Table 2: The corresponding labels of the category numbers in Fig. 5 b.
Dataset No. Label
BreastDM 0 Benign
1 Malignant
ADNI 0 Cognitively Normal (CN)
1 Mild Cognitive Impairment (MCI)
2 Alzheimer’s Disease (AD)
LLD-MMRI 0 Hepatocellular Carcinoma (HCC)
1 Intrahepatic Cholangio Carcinoma (ICC)
2 Hepatic Metastasis (HM)
3 Hepatic Cyst (HC)
4 Hepatic Hemangioma (HH)
5 Focal Nodular Hyperplasia (FNH)
6 Hepatic Abscess (HA)
LUNA16 0 Non-nodule
1 Nodule
OrganMNIST3D 0 Heart
1 Left Lung
2 Right Lung
3 Liver
4 Spleen
5 Pancreas
6 Left Kidney
7 Right Kidney
8 Bladder
9 Left Femoral Head
10 Right Femoral Head
Table 3: Parameter setting in the pre-training phase.
Encoder Scale Parameter Value
nnUNet - Learning rate 0.0001
Swin-Transformer Base Learning rate 0.000001
Feature size 48
Bottleneck Depth 768
Large Learning rate 0.000001
Feature size 96
Bottleneck Depth 1,536
Huge Learning rate 0.000001
Feature size 192
Bottleneck Depth 3,072
Common parameter Value
Optimizer AdamW
Number step 200,000
Warmup step 1,000
Learning rate schedule Cosine
Batch size 8
Roi x,y,z 96
Figure 7: Visual representation of the demographics of Triad-131K and its three subsets in this study.
Figure 8: Decoder architecture used uniformly in the pre-training phase.
Acknowledgements
This research is supported in part by the National Institutes of Health under Award Number R56EB033332, R01EB032680, R01DE033512 and R01CA272991.
Contributions
Shansong Wang: Writing-original draft, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Mojtaba Safari: Data curation, Writing-review & editing. Qiang Li: Investigation, Writing-review & editing. Chih-Wei Chang: Writing-review & editing. Richard LJ Qiu: Writing-review & editing. Justin Roper: Writing-review & editing. David S. Yu: Writing-review & editing. Xiaofeng Yang: Writing-review & editing, Supervision, Resources, Project administration, Funding acquisition.